OpenStudy (anonymous):

Could I fit a 6' piece of lumber in my trunk and close it. My trunk measures 3'x4'x2'

jimthompson5910 (jim_thompson5910):

anyways, imagine a 3D rectangular prism like this 

jimthompson5910 (jim_thompson5910):

the space diagonal would stretch from one corner to the opposite corner  that is effectively the longest straight line you can draw inside the prism

jimthompson5910 (jim_thompson5910):

you use the 3D version of the pythagorean theorem to find this length \[\Large d = \sqrt{a^2 + b^2 + c^2}\] \[\Large d = \sqrt{3^2 + 4^2 + 2^2}\] \[\Large d = \sqrt{29}\] \[\Large d \approx 5.3851648\] The diagonal's length is approximately 5.3851648 feet long. So this is the upper limit. You'll have less than this because the board isn't infinitely thin. The thickness or width of the board will make sure you have less room to work with.

OpenStudy (kendricklamar2014):

Is this school work?? If it's not, get your tape measure and measure 6' on your trunk and see if it fits